1. what is the circumference of a circle with a diameter of 30 centimeters?
2. for a botton to fit through its bottonhole, the hole needs to be the size of the botton's diameter. what size bottonhole is needed for a botton with a circumference of 9.42 centimeters?
3. The spoke of a wheel reaches from the center of the wheel to its rim. if the circumference of the rim of the wheel is 42 inches, how long is each spoke? round to the nearest hundreth.

Answer by checkley71(8403)

C=2PIR
C=2*3.14*15
C=94.2 CM. FOR THE CIRCUMFERANCE.
--------------------------------------
9.42=PID
9.42=3.14D
D=9.42/3.14
D=3 CM.
----------------------------------------
42=2PIR
42=2*3.14*R
42=6.28R
R=42/6.28
R=6.69 CM.IS THE LENGTH OF THE SPOKES.